SmartBeans is Canada’s largest online retailer of fair-trade organic coffee, and imports coffee beans from several countries. SmartBeans operates 300 days a year and sells an average of 150 pounds of Harar (Ethiopia) Fair Trade Organic beans a day. After ordering, beans are always shipped from Ethiopia within exactly 11 days. Annual holding costs per pound are estimated to be 15% of the per pound cost of beans. The ordering cost is $35 per order. The cost of a pound of green coffee beans is $7.

What is the economic order quantity (EOQ) for Harar Fair Trade Organic coffee beans? What is the average inventory? What is the expected total number of orders per year?
Answer #1

Given Annual demand D = 150*300 = 45000

Cost = $7

Holding cost H = 15% of the per pound cost of beans = 0.15*7 = $1.05

Ordering cost S = 35

Lead time L = 11 days

EOQ=\sqrt{\frac{2DS}{H}}

EOQ=\sqrt{\frac{2*45000*35}{1.05}}

EOQ = 1732

Average inventory = EOQ/2 = 1732/2 = 866

Expected total number of orders per year = (D/EOQ) = 45000/1732 = 26